OpenStudy (anonymous):

can someone help me? Factor: x^2 - 13x + 36

hero (hero):

Hint: x^2 - 4x - 9x + 36

OpenStudy (anonymous):

x^2-9x-4x+36

OpenStudy (anonymous):

x^2-13x+36

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

then i wrote -13x+36=-x^2 36=-x^2+13x 36/13=-^2+13x/13 2=x^3 ? <<<is that right?

hero (hero):

Not exactly.

hero (hero):

Here's another hint: x^2 - 4x - 9x + 36 x(x - 4) - 9(x - 4)

OpenStudy (anonymous):

(x-4)(x-9)?

hero (hero):

Yes, finally...

hero (hero):

You can check it by multiplying it back out: (x - 4)(x - 9) = x(x - 9) - 4(x - 9) = x^2 - 9x - 4x + 36 = x^2 - 13x + 36